# E Meena Nanivadekar A Azad Shenoy Roy Chowdhuri Banga Chhabra Bhat Kalluri Chetupalli Vijayasenan Ganapathy

### 04 Mar 2026 v2

@@ Abstract
The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge introduces a conversational AI benchmark focused on understanding goal-oriented, real-world medical dialogues collected in the field. The challenge addresses multi-speaker interactions between healthcare workers and seekers, characterized by spontaneous, noisy and overlapping speech across Indian languages and dialects. As part of the challenge, medical conversational dataset comprising 25 hours of development data and 10 hours of blind evaluation recordings was released. We provided baseline systems within a unified end-to-end pipeline across 4 tasks - speaker diarization, automatic speech recognition, topic identification and dialogue summarization - to enable consistent benchmarking. System performance is evaluated using established metrics such as diarization error rate (DER), time-constrained minimum-permutation word error rate (tcpWER), and ROUGE-L. During this evaluation (Phase-I), $12$ teams, across the globe, actively participated pushing the baseline systems on these metrics. However, even with a 686-8 week dedicated effort from various participants, the task is shown to be substantially challenging, and the existing systems are significantly short of healthcare deployment readiness.
@@ 1 Introduction
Speech processing tools for the understanding of conversations from the healthcare domain can have transformative effects on public health systems, especially in community healthcare settings. Existing speech datasets in this domain are mostly recorded in controlled environments such as hospitals, involve structured clinician-patient conversations, and are predominantly in English. AI-driven tools developed using such datasets often fail to perform when deployed in real-world setups involving community healthcare workers. Further, there is a lack of research in low-resource languages, in particular in Indian languages. To catalyze research efforts in this space, we collected and released an annotated corpus of conversations between community health workers and healthcare seekers in informal settings in Hindi. Using the data, we hosted an open leaderboard-style challenge - ``The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge''.
DISPLACE-M makes three main contributions:
@@
Figure 1: Timeline of DISPLACE-M
2 Related Work
ASR research has seen significant progress in conversational settings, particularly for meetings, call centers, and other multi-party environments. Datasets such as the ICSI Meeting Corpus [ICSI_Meeting_dataset] and the AMI Meeting Corpus [ami_datasset] have been widely used to study spontaneous multi-speaker conversational speech. However, the conversations in these datasets differ significantly from those in frontline healthcare settings, where conversations are goal-driven, medically grounded, and often involve nuanced contextual information. Topic classification has been widely studied for meetings, call centers, and online discussions, while dialogue summarization datasets have largely been derived from customer support conversations, news interviews, or scripted meeting scenarios [gliwa-etal-2019-samsum, Zhong2021QMSumAN].
@@ 2 Related Work
In the Indian context, there has been a growing interest in developing ASR systems for low-resource and multilingual settings. However, existing datasets are restricted to isolated utterances or short conversational clips [Bhogale2022EffectivenessOM, bhogale23_interspeech, Nithya2023SPRINGINXAM, javed-etal-2024-indicvoices], which limits their suitability for studying long-form, dialogue centric ASR systems. Further, there are no public datasets covering conversations in the healthcare domain.
Despite substantial progress in medical NLP and conversational ASR, there is still no benchmark that captures the complexity of real-world frontline healthcare interactions. The DISPLACE-M Challenge addresses these limitations by introducing a multi-task benchmark spanning diarization, ASR, topic identification, and dialogue summarization.
@@
The Phase-I of the DISPLACE-M evaluations was organized between January 1, 2026 and February 27, 2026. Timeline is shown in Figure 1 . The details of the dataset (Section 3), tracks and metrics (Section 4), baseline systems (Section 5) and challenge results (Section 6) are described below.
3 DISPLACE-M Corpus
@@ 3.1 Data Collection
We recorded around $50$ hours of annotated healthcare conversational speech from rural and semi-urban areas, collected through field deployments in the Indian states of Haryana and Bihar. Data collection was conducted by $80$ frontline community health workers across $20$ administrative blocks spanning $10$ districts. The primary participants were NPHWs (non-physician health workers) - ASHA [^1] ASHA (Accredited Social Health Activists) workers, who are government-appointed female community health workers serving as the first point of contact for healthcare seekers (HS) and Anganwadi Sevikas  [^2] Anganwadi sevikas are frontline workers under India’s Integrated Child Development Services (ICDS) program. Recordings captured natural NPHW-participant conversations occurring during routine community healthcare interactions. These interactions took place in real care environments such as Primary Health Centres (PHCs), Anganwadi centers, primary schools, open village meeting spaces (chaupals), household visits and community outreach settings. In our recordings, Hindi served as the primary language with code-switching to Indian English and with heavy use of regional dialects, including Haryanvi, Bhojpuri, and Magahi, reflecting natural linguistic practices.
The recorded conversations include speakers from different age groups. Each recording ranges from 55-$30$ minutes, typically involving goal-oriented conversations between two participants - NPHW and a healthcare seeker. The conversations, which include nature field ambiance, primarily focused on topics such as General Health and Constitutional Symptoms (e.g., weakness, fatigue, weight loss, low hemoglobin/blood loss), Women’s Health and Gynecological Issues (e.g., menstrual problems, pregnancy, family planning), Acute Illnesses and Infections (e.g., fever, cough, cold, typhoid, dengue), and Preventive Care and Diagnostics (e.g., routine check-ups, blood tests, ultrasounds). All audio recordings were captured using the built-in microphone of a mobile device. The recording format was WAV using $16$-bit PCM at 1616kHz sampling rate.
@@ 3.2 Annotation
All recordings were processed through a structured, multi-stage annotation pipeline prior to dataset finalization. We performed a multi-stage check to ensure quality of the audio files, accuracy of the time-markers and the correctness of the ASR transcripts. Further, the clinical summary is given by an expert clinician (Doctor) who had access to the audio and ground-truth rich transcriptions. The workflow was organized into distinct segmentation, transcription, and quality assurance phases to ensure traceability and reproducibility. We include the details of the dataset, pre-processing steps, and quality-checks in the supplementary material.
@@ 4 Challenge Tracks and Metrics
    Track 2 - Automatic Speech Recognition (ASR) system transcribes multi-speaker, mono-lingual (Hindi) and code-mixed (with English words) healthcare conversations. Systems are expected to generate time-marked word-level transcriptions for the spontaneous conversational speech. The performance metric is character error rate (CER), word-error rate (WER) and tcpWER (Time-Constrained minimum-Permutation Word Error Rate) [neumann23_chime]. The tcpWER extends standard WER to multi-speaker settings by computing the minimum WER over all possible permutations of reference and hypothesized speaker streams after concatenating each speaker’s utterances [neumann23_chime] with constraints along the temporal axis for alignment. It therefore jointly reflects transcription quality and speaker assignment consistency, making it suitable for evaluating end-to-end diarization-aware ASR.
@@ 5 Baseline Systems
For the baseline system development, we follow a cascaded system approach, illustrated in Fig. 2. Given raw conversational audio, we first perform speaker diarization to obtain time-stamped speaker segments. These segments are subsequently used to guide the ASR module, which generates speaker-attributed transcripts. The resulting transcripts then serve as inputs to topic identification and dialogue summarization.
@@ 5.1 Track-1: Speaker Diarization (SD)
For the DISPLACE-M baseline, DiariZen was evaluated under two setups - zero-shot inference and supervised fine-tuning with the development data. The fine-tuning was done with the AHC threshold set to $0.7$ and minimum cluster size set to $30$. This system is referred to as Baseline-2.
@@ 5.3 Track-3: Topic Identification (TI)
We employ an ASR–LLM cascade pipeline to identify medical topics in healthcare conversations. This approach first converts spontaneous speech into text using ASR systems, followed by an LLM, medgemma-1.5-4b-it [sellergren2025medgemma] prompting to generate the relevant health topics.
@@ 5.4 Track 4: Dialogue Summarization (DS)
Similar to the topic identification pipeline, we used the ASR transcript for the dialogue summarization (DS) task. For the summarization stage, we used LLAMA-3.2-3B model [mansha2025resource] as the text-based summarization model in a zero-shot setting. The prompt used for summarization is described in the supplementary material.
@@ 6 Challenge Results
We also benchmarked state-of-the-art closed-source models as reference systems. In particular, we evaluated Gemini 2.5 Pro [comanici2025gemini] across all four tracks and Sarvam AI’s Saaras v3777https://dashboard.sarvam.ai/speech-to-text for Track 1 (Speaker Diarization) and Track 2 (Automatic Speech Recognition).
@@ 6.1 Track 1: Speaker Diarization
Figure 3 illustrates the Eval Phase-I results for the SD track. A total of $9$ teams participated in Phase-I evaluation. The top four systems achieved moderate improvements over Baseline-2, as indicated by the green bars in the figure.
Figure 3: Comparison of Diarization Error Rate (DER %) for Track 1 Evaluation Phase-I across submitted teams.
T1 used a hybrid end-to-end system trained on multiple datasets and fine-tuned it on the development sets. They also employed a Dynamic Logits Fusion strategy to combine 5 complementary systems, and obtained their best results. T2 and T3 fine-tuned the large variant of DiariZen with the WavLM model. T2 used Dev-1 for training and Dev-2 for validation, and obtained their best results with learning rates of $10^{-3}$ and $2\times 10^{-5}$ for the conformer blocks and WavLM, respectively. In contrast, T3 used a mix of both the development sets, employing a $80:20$ train-validation split. They utilize a three-stage progressive fine-tuning strategy with different hyperparameter values in each stage.
@@ 6.2 Track 2: ASR
The ASR track (Track-2) received $4$ submissions, of which two teams outperformed the baseline system. Table 1 presents the Eval Phase-I results for the ASR track. Among the baselines described in Section 5.2, fine-tuning significantly improved performance over the zero-shot setting. The fine-tuned IndicConformer achieved a tcpWER of 20.2320.23% compared to 26.7826.78% in the zero-shot setup, indicating the importance of in-domain adaptation.
@@
Table 1: ASR Track (Track-2) Results on Eval Phase-I; ZS = Zero-shot inference, FT = Fine-tuned on challenge dev-set.
@@ 6.3 Track 3: Topic Identification
Table 2 presents the results for both Track $3$ and Track $4$. Track $3$ received two submissions in Eval Phase-I and both teams outperformed the baseline system described in Sec. 5.4.
@@
Team 1 (T1) achieved the highest ROUGE scores, obtaining a ROUGE-1 of $0.46$ and a ROUGE-L of $0.44$, by using the Gemini 3 Pro end-to-end model operating directly on raw audio in a zero-shot configuration. Team 2 (T2) enhanced the baseline ASR-LLM pipeline by improving translation, topic extraction and incorporating auxiliary patient information. They included supplementary cues from raw speech, such as vocal characteristics, to refine topic prediction; for instance, knowledge of a patient’s sex may help distinguish between stomach pain and menstrual pain. They replaced the baseline translation model with Llama 3.1 70B Instruct and increased the decoding length to 2048 to better capture conversational context. For topic extraction, they employed Llama3.1-Aloe-Beta-8B, a model fine-tuned for medical reasoning with prompts refined iteratively and guided by a reference list of development-set topics while allowing open-set outputs. Additionally, diarization outputs were used to isolate patient speech, from which age and biological sex were estimated using a Wav2Vec2-based model. These attributes were appended to the prompt to enable age- and sex-aware topic identification. Their approach relies entirely on pre-trained models without additional fine-tuning, using only the released development sets for validation.
@@ 6.4 Track 4: Dialogue Summarization
Track $4$ received three submissions in Eval Phase-I, of which two teams outperformed the baseline system. The top-performing team (T1) proposed an end-to-end pipeline combining Silero-VAD [Silero_VAD] for segmentation, speech translation from Hindi to English using AudioX-North Whisper [jiviai_audiox_north_2024] and GPT-4o-mini-based summarization using a structured clinical prompt. Team 2 (T2) used the Gemini-3 Pro end-to-end model in a $6$-shot configuration, where raw audio was directly provided as input without an intermediate ASR step and the model was guided by six example transcript–summary pairs from the challenge development set.
@@ 7 Discussion and Conclusion
The DISPLACE-M challenge aims to study and investigate the difficulties that state-of-the-art systems face in transcribing and understanding frontline health conversations. Our results indicate that upstream tasks such as speaker diarization and automatic speech recognition still requires work to enable seamless understanding of these conversations. The best ASR system achieves a WER of $18.2$ % (and tcpWER of $18.65$ %). Further, downstream tasks such as topic identification and dialogue summarization remain significantly more challenging due to their domain-specific nature and the complexity of interactions. Among the evaluated tasks, dialogue summarization is the most challenging and we observe that even large closed-source models struggle to generate reliable and clinically accurate medical summaries. One key reason is the conversational structure of health worker-seeker interactions, as they often contain implicit symptoms, fragmented descriptions and context-dependent dialogue, which require deeper reasoning to interpret correctly. Generating accurate summaries requires domain knowledge and multi-step reasoning to interpret symptoms and identify medically relevant information. Since the timeline was only about $6$ weeks, we have initiated a Phase-II plan for this evaluation which allows $10$ more weeks of open-leaderboard style evaluation. The Phase-II evaluation will include more languages (beyond Hindi) and will likely attract more participation.

### 05 Mar 2026 v3

@@ Abstract
The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge introduces a conversational AI benchmark for understanding goal-oriented, real-world medical dialogues. The challenge addresses multi-speaker interactions between frontline health workers and care seekers, characterized by spontaneous, noisy and overlapping speech. As part of the challenge, medical conversational dataset comprising $40$ hours of development and $15$ hours of blind evaluation recordings was released. We provided baseline systems across 4 tasks - speaker diarization, automatic speech recognition, topic identification and dialogue summarization - to enable consistent benchmarking. System performance is evaluated using diarization error rate (DER), time-constrained minimum-permutation word error rate (tcpWER) and ROUGE-L. This paper describes the Phase-I evaluation - data, tasks and baseline systems - along with the summary of the evaluation results.
@@ 1 Introduction
Speech processing tools for the understanding of conversations from the healthcare domain can have transformative effects on public health systems, especially in community healthcare settings. Existing speech datasets in this domain are mostly recorded in controlled environments such as hospitals, involve structured clinician-patient conversations, and are predominantly in English. AI-driven tools developed using such datasets often fail to perform when deployed in real-world setups involving community healthcare workers. Further, there is a lack of research in multi-lingual settings. To catalyze research efforts in this space, we collected and released an annotated corpus of conversations between community health workers and healthcare seekers in informal settings in Hindi. Using the data, we hosted an open leaderboard-style challenge - ``The DIarization and Speech Processing for LAnguage understanding in Conversational Environments - Medical (DISPLACE-M) challenge''. This paper makes three main contributions:
@@
2 Related Work and Paper Organization
ASR research has seen significant progress in conversational settings, particularly for meetings, call centers, and other multi-party environments. Datasets such as the ICSI Meeting Corpus [ICSI_Meeting_dataset] and the AMI Meeting Corpus [ami_datasset] have been widely used to study spontaneous multi-speaker conversational speech. However, the conversations in these datasets differ significantly from those in frontline healthcare settings, where conversations are goal-driven, medically grounded, and often involve nuanced contextual information. Topic classification has been widely studied for meetings, call centers, and online discussions, while dialogue summarization datasets have largely been derived from customer support conversations, news interviews, scripted meeting scenarios or informal meetings [gliwa-etal-2019-samsum, Zhong2021QMSumAN, baghel2024summary, kalluri2024second].
@@ 2 Related Work and Paper Organization
In the Indian context, there has been a growing interest in developing diarization and ASR systems for multilingual settings [baghel2023displace]. However, existing datasets are restricted to isolated utterances or short conversational clips [Bhogale2022EffectivenessOM, bhogale23_interspeech, Nithya2023SPRINGINXAM, javed-etal-2024-indicvoices], which limits their suitability for studying long-form, dialogue centric ASR systems. Further, there are no public datasets covering conversations in the healthcare domain. Despite substantial progress in medical NLP and conversational ASR, there is still no benchmark that captures the complexity of real-world frontline healthcare interactions.
@@
To fill this gap, we collected the DISPLACE-M dataset and launched the evaluation campaign. The Phase-I of the DISPLACE-M evaluations was organized between January 1, 2026 and February 27, 2026. The details of the dataset (Section 3), tracks and metrics (Section 4), baseline systems (Section 5) and challenge results (Section 6) are described below.
3 DISPLACE-M Challenge
@@ 3.1 Data Collection
We recorded around $55$ hours of annotated healthcare conversational speech from rural and semi-urban areas, collected through field deployments in the Indian states of Haryana and Bihar. Data collection was conducted by $80$ frontline community health workers across $20$ administrative blocks spanning $10$ districts. The primary participants were NPHWs (non-physician health workers) - ASHA [^1] ASHA (Accredited Social Health Activists) workers, who are government-appointed female community health workers serving as the first point of contact for healthcare seekers (HS) and Anganwadi Sevikas  [^2] Anganwadi sevikas are frontline workers under India’s Integrated Child Development Services (ICDS) program and healthcare seekers (HS). Recordings captured natural NPHW-HS conversations occurring during routine community healthcare interactions. These interactions took place in real care environments such as Primary Health Centres (PHCs), Anganwadi centers, primary schools, open village meeting spaces (chaupals), household visits and community outreach settings. In our recordings, Hindi served as the primary language with code-switching to Indian English and with heavy use of regional dialects, such as Haryanvi, Bhojpuri, and Magahi, reflecting natural linguistic practices.
The recorded conversations include speakers from different age groups. Each recording ranges from 55-$30$ minutes, typically involving goal-oriented conversations between two participants - NPHW and a healthcare seeker. The conversations, which include nature field ambiance, primarily focused on topics such as General Health and Constitutional Symptoms (e.g., weakness, fatigue, weight loss, low hemoglobin/blood loss), Women’s Health and Gynecological Issues (e.g., menstrual problems, pregnancy, family planning), Acute Illnesses and Infections (e.g., fever, cough, cold, typhoid, dengue), and Preventive Care and Diagnostics (e.g., routine check-ups, blood tests, ultrasounds). All audio recordings were captured using the far-field microphone of a mobile device. The recording format was WAV using $16$-bit PCM at 1616kHz sampling rate.
@@ 3.2 Annotation
All recordings were processed through a structured, multi-stage manual annotation pipeline prior to dataset finalization. We performed a multi-stage check to ensure quality of the audio files, accuracy of the time-markers and the correctness of the ASR transcripts. Further, the clinical summary is given by an expert clinician (Doctor) who had access to the audio and ground-truth rich transcriptions. The workflow was organized into distinct segmentation, transcription, and quality assurance phases to ensure traceability and reproducibility. More details of the dataset, pre-processing steps, and the annotation quality-checks are provided in the supplementary material.
@@ 4 Challenge Tracks and Metrics
    Track 2 - Automatic Speech Recognition (ASR) system transcribes multi-speaker healthcare conversations. Systems are expected to generate time-marked word-level transcriptions for the spontaneous conversational speech. The performance metric is character error rate (CER), word-error rate (WER) and tcpWER (Time-Constrained minimum-Permutation Word Error Rate) [neumann23_chime]. The tcpWER extends standard WER to multi-speaker settings by computing the minimum WER over all possible permutations of reference and hypothesized speaker streams after concatenating each speaker’s utterances [neumann23_chime] with constraints along the temporal axis for alignment. It therefore jointly reflects transcription quality and speaker assignment consistency, making it suitable for evaluating end-to-end diarization-aware ASR.
@@ 5 Baseline Systems
For the baseline system development, we follow a cascaded system approach illustrated in Figure 1. Given raw conversational audio, we first perform speaker diarization to obtain time-stamped speaker segments. These segments are subsequently used to guide the ASR module, which generates speaker-attributed transcripts. These transcripts then serve as inputs to topic identification and dialogue summarization.
@@ 5.1 Track-1: Speaker Diarization (SD)
For the DISPLACE-M baseline, DiariZen was evaluated under two setups - zero-shot inference (Baseline-1) and supervised fine-tuning with the development data (Baseline-2). The fine-tuning was done with the AHC threshold set to $0.7$ and minimum cluster size set to $30$.
@@ 5.3 Track-3: Topic Identification (TI)
We employ an ASR–LLM cascade pipeline to identify medical topics in healthcare conversations. This approach first converts spontaneous speech into text using ASR systems, followed by an LLM - medgemma-1.5-4b-it [sellergren2025medgemma] - to generate the relevant health topics. The specific prompt and setting used is described in Appendix A.5.1
@@ 5.4 Track 4: Dialogue Summarization (DS)
Similar to the topic identification pipeline, we used the ASR transcript for the dialogue summarization (DS) task. For the summarization stage, we used LLAMA-3.2-3B model [mansha2025resource] as the text-based summarization model in a zero-shot setting. The prompt used for summarization is described in Appendix A.5.2
@@ 6 Challenge Results
We also benchmarked state-of-the-art closed-source models as reference systems. In particular, we evaluated Gemini 2.5 Pro [comanici2025gemini] across all four tracks and Sarvam AI’s Saaras v3777https://dashboard.sarvam.ai/speech-to-text for Track 1 (Speaker Diarization) and Track 2 (ASR).
@@ 6.1 Track 1: Speaker Diarization
Figure 2 illustrates the Eval Phase-I results for the SD track. A total of $9$ teams participated in Phase-I evaluation. The top four systems achieved improvements over Baseline-2,.
Figure 2: Comparison of Diarization Error Rate (DER %) for Track 1 Evaluation Phase-I DS task.
Team 1 (T1) used a hybrid end-to-end system trained on multiple datasets and fine-tuned it on the development sets. They also employed a Dynamic Logits Fusion strategy to combine 5 complementary systems, and obtained their best results. Team 2 (T2) and Team 3 (T3) fine-tuned the large variant of DiariZen with the WavLM model. T2 used Dev-1 for training and Dev-2 for validation, and obtained their best results with learning rates of $10^{-3}$ and $2\times 10^{-5}$ for the conformer blocks and WavLM, respectively. In contrast, T3 used a mix of both the development sets, employing a $80:20$ train-validation split. They utilize a three-stage progressive fine-tuning strategy with different hyperparameter values in each stage.
@@ 6.2 Track 2: ASR
The ASR track received $4$ submissions, of which two teams outperformed the baseline system. Table 1 presents the Eval Phase-I results. Among the baselines described in Section 5.2, fine-tuning significantly improved performance over the zero-shot setting. The fine-tuned IndicConformer achieved a tcpWER of 20.2320.23% compared to 26.7826.78% in the zero-shot setup, indicating the importance of in-domain adaptation.
@@
Table 1: Results on Track-2 (ASR); ZS = Zero-shot inference, FT = Fine-tuned on challenge dev-set.
@@ 6.3 Track 3: Topic Identification
Table 2 presents the results for both Track $3$ and Track $4$. Track $3$ received two submissions in Eval Phase-I and both teams outperformed the baseline system described in Section 5.4.
@@
Team 1 (T1) achieved the highest ROUGE scores, obtaining a ROUGE-1 of $0.46$ and a ROUGE-L of $0.44$, by using the Gemini 3 Pro end-to-end model operating directly on raw audio in a zero-shot configuration. Team 2 (T2) enhanced the baseline ASR-LLM pipeline by improving translation, topic extraction and incorporating auxiliary patient information. They included supplementary cues from raw speech, such as vocal characteristics, to refine topic prediction; for instance, knowledge of a patient’s sex may help distinguish between stomach pain and menstrual pain. They replaced the baseline translation model with Llama 3.1 70B Instruct and increased the decoding length to 2048 to better capture conversational context. For topic extraction, they employed Llama3.1-Aloe-Beta-8B, a model fine-tuned for medical reasoning with prompts refined iteratively and guided by a reference list of development-set topics. Additionally, diarization outputs were used to isolate patient speech, from which age and biological sex were estimated using a Wav2Vec2-based model. These attributes were appended to the prompt to enable age- and sex-aware topic identification. Their approach relies entirely on pre-trained models without additional fine-tuning, using only the released development sets for validation.
@@ 6.4 Track 4: Dialogue Summarization
Track $4$ received three submissions in Eval Phase-I, of which two teams outperformed the baseline system. The top-performing team (T1) proposed an end-to-end pipeline combining Silero-VAD [Silero_VAD] for segmentation, speech translation from Hindi to English using AudioX-North Whisper [jiviai_audiox_north_2024], and GPT-4o-mini-based summarization using a structured clinical prompt. Team 2 (T2) used the Gemini-3 Pro end-to-end model in a $6$-shot configuration, where raw audio was directly provided as input without an intermediate ASR step and the model was guided by six example transcript–summary pairs from the challenge development set.
@@ 7 Discussion and Conclusion
The DISPLACE-M challenge aims to study and investigate the difficulties that state-of-the-art systems face in transcribing and understanding frontline health conversations. Our results indicate that upstream tasks such as speaker diarization and automatic speech recognition still require improvements to enable seamless understanding of these conversations. The best ASR system achieved a WER of $18.2$ % (and tcpWER of $18.65$ %). Further, downstream tasks such as topic identification and dialogue summarization remain significantly more challenging due to their domain-specific nature and the complexity of interactions. Among the evaluated tasks, dialogue summarization is found to be the most challenging and we observe that even large closed-source models struggle to generate reliable and clinically accurate medical summaries. One key reason is the conversational structure of health worker-seeker interactions, as they often contain implicit symptoms, fragmented descriptions and context-dependent dialogue, which require deeper reasoning to interpret correctly. Generating accurate summaries requires domain knowledge and multi-step reasoning to interpret symptoms and identify medically relevant information. Since the timeline was only about $6$ weeks, we have initiated a Phase-II plan for this evaluation which allows $10$ more weeks of open-leader-board style evaluation. The Phase-II evaluation will include more languages (beyond Hindi).